### v1

# Neural Spelling: A Spell-Based BCI System for Language Neural Decoding

Abstract
Brain-computer interfaces (BCIs) present a promising avenue by translating neural activity directly into text, eliminating the need for physical actions. However, existing non-invasive BCI systems have not successfully covered the entire alphabet, limiting their practicality. In this paper, we propose a novel non-invasive EEG-based BCI system with Curriculum-based Neural Spelling Framework, which recognizes all 26 alphabet letters by decoding neural signals associated with handwriting first, and then apply a Generative AI (GenAI) to enhance spell-based neural language decoding tasks. Our approach combines the ease of handwriting with the accessibility of EEG technology, utilizing advanced neural decoding algorithms and pre-trained large language models (LLMs) to translate EEG patterns into text with high accuracy. This system show how GenAI can improve the performance of typical spelling-based neural language decoding task, and addresses the limitations of previous methods, offering a scalable and user-friendly solution for individuals with communication impairments, thereby enhancing inclusive communication options.
Index Terms:
Fuzzy Logic, Transformers, fNIRS, Social Neuroscience
I Introduction
Brain-computer interfaces (BCIs) have emerged as a pivotal area of research within human-computer interaction (HCI), distinguished by their capacity to seamlessly integrate neural signals with external systems. Pioneering studies such as those by Guo et al.[1], Chen et al.[2], Cao et al.[3], and Lin et al.[4] underscore BCIs’ role in advancing neuroscience and technology. These interfaces create direct communication pathways that are especially beneficial for individuals with limited speech or motor functions. Language decoding represents a critical domain within BCI research, aimed at deciphering the neural correlates of speech and language processing. This capability not only enhances interaction paradigms but also opens new avenues for communication, offering significant improvements in quality of life for those with severe communicative impairments.
Figure 1: Demonstrations of two typical language decoding frameworks. (A) Direct speech synthesis approach. (B) Spelling-based approach using phonemes. (C) Direct text synthesis approach. (D) Spelling-based approach using letters.
Initially, BCI research focused on visual-based decoding approaches, such as steady-state visually evoked potentials (SSVEP)[1], which, despite their reliability, demand high cognitive effort and are unsuitable for prolonged use[2]. These methods often fail to align with natural human language patterns, posing significant usability challenges. The advent of invasive neural decoding technologies marked a significant advancement, allowing for direct interpretation of brain signals via electrocorticography (ECoG) or stereo-EEG (sEEG)[5, 6]. These methods have demonstrated substantial improvements in user performance, significantly enhancing communicative capacities for patients with speech impairments. However, the invasive nature, high cost, and ethical concerns limit their general applicability[7]. In contrast, non-invasive BCIs, predominantly utilizing electroencephalography (EEG), offer a more accessible alternative. These systems are less obtrusive and more cost-effective, broadening potential user demographics [8]. Despite the challenges of signal noise and the extensive training required for users, recent studies have demonstrated EEG’s potential in effective language decoding [9, 10].
With the rise of Generative AI (GenAI), the integration of large language models (LLMs) into BCI research has opened new avenues for enhancing language decoding [11]. This integration can be implemented within two distinct frameworks: directly synthesizing speech or text from brain signals using pre-trained LLMs, as illustrated in Fig. 1(A) and Fig. 1(C) respectively; and decoding brain signals into minimal language units such as phonemes, as shown in Fig. 1(B), or letters, as depicted in Fig. 1(D), followed by text generation through natural language processing (NLP) models.
Figure 2: (A): The Experiment Design. (B): The trajectory finishing with time. (C): The PSD feature calculated from Raw EEG signal. (D): The structures of the Trajectory Resnet18 Encoder and the Converlutional EEG Encoder. (E): The Letter Probability Distribution in the outputs from the classifier layers. (F): The structure of the sentence generator.
The first framework often faces limitations due to the disparity in data scales between brain data and text/image datasets, which can restrict the system to operating within predefined datasets, merely retrieving sentences rather than generating novel content. The second approach, known as the spell-based method [12, 13], operates by initially decoding neural signals into their minimal representational units, such as the 26 letters, or even nine-key input (T9) [14]. It subsequently employs GenAI to construct coherent text in a second stage. This technique is advantageous as it requires fewer categories for the neural decoding model, thereby reducing both the cognitive load on the user and the difficulty associated with model training. A significant benefit of this method is its utilization of extensive textual datasets in training GenAI, particularly for generative error correction (GEC) [15], which enhances the accuracy and flexibility of the generated text. Moreover, this method is not constrained by the size of the neural training samples and can adapt swiftly to new linguistic content through fine-tuning of the GenAI, circumventing the need for extensive retraining on new brain data.
[…]
This paper contributes to the evolving landscape of brain-computer interfaces (BCIs) by proposing a hybrid approach that merges the accessibility of non-invasive EEG with the advanced linguistic capabilities of GenAI. We introduce a novel curriculum-based neural spelling framework (CNS) for neural language decoding. This framework employs a convolutional neural network (CNN)-based encoder to initially learn individual-specific letter transition patterns, followed by a curriculum-driven LLM to synthesize sentence texts. The results demonstrate significant enhancements in the performance of language decoding tasks. By integrating these advanced technologies, we aim to facilitate more natural, efficient, and inclusive communication methods for individuals with varying physical abilities, heralding a new era in assistive communication technologies. This introduction outlines the transformation of BCI technology from basic visual decoding to sophisticated language synthesis, setting the stage for a detailed discussion on how current technologies can be integrated and enhanced to better meet the communicative needs of a diverse user population.
The contributions of this work are threefold:
  • This study is the first to collect and analyze brain dynamics patterns related to EEG-based handwriting of all 26 letters.
  • We employ a CNN-based model for EEG encoding, which achieves exemplary top-k accuracy, averaging across all subjects.
  • We propose a novel curriculum supervised fine-tuning method that enables an LLM to learn subject-specific letter transition patterns effectively, thereby enabling high-quality sentence synthesis. This approach has shown promising results, achieving high scores on established metrics.
II Related Work
II-A Language Decoding
Recent advances in Language Neural Decoding have significantly enhanced BCI systems. RNN-based models have achieved a $23.8\%$ word error rate on a 125,000-word vocabulary using chronic ECoG signals, highlighting their potential as speech neuroprostheses [7]. ECoG-based Speech BCIs have demonstrated stable control of assistive devices for up to three months with minimal calibration, supporting daily unassisted use [19]. Additionally, speech synthesis from ECoG signals enables decoding spoken sentences [5], while contrastive learning models have decoded perceived speech from non-invasive recordings like EEG and MEG, enriching decoding techniques [9]. To better predict the text, a novel task, named Cross-Modal Cloze (CMC) task [20], which is to predict the target word with a context as prompt, are proposed, achieving $28.91\%$ accuracy.
[…]
ECoG systems have provided communication solutions for late-stage ALS patients by decoding English letters [22], while silent spelling mechanisms predict characters in 2.5 seconds, improving practical use [23]. CNNs have decoded sentences [24] and simple commands [25, 26], while EEG-based user authentication systems using dynamic signatures have been explored for security [27, 28].
II-B Enhancing Text-to-Text Error Correction with Large Language Models
Advances in NLP have been fueled by the emergence of LLMs, which represent a paradigm shift from traditional supervised learning to pretrained models that leverage vast corpora, subsequently fine-tuned or prompted for specific tasks [29, 30]. These models, primarily encompassing BERT-based architectures with bi-directional attention mechanisms [31] and GPT-based models utilizing auto-regressive learning schemas [32, 33], excel at building contextual representations and deriving nuanced understanding from extensive natural language data. As a result, complex language tasks such as document-level translation [34, 35] and advanced question answering [36] can now be performed with minimal supervision.
When it comes to GEC area, LLMs excel at text manipulation tasks such as grammar correction [37], text rewriting [38], and improving spoken language comprehension [39, 40]. By integrating generation and re-ranking, LLMs provide dynamic solutions for error correction and text refinement [41, 42]. Models like ChatGPT [43] and ChatGLM [44] expand these capabilities further, performing text understanding, generation, and correction in unified frameworks. Despite their computational demands, smaller models like BART [45] and T5 [46] offer resource-efficient alternatives for specialized tasks, maintaining robust performance in constrained environments [47].
By integrating LLM-based generative error correction, we can re-organize the slightly disordered output sequence of the neural spelling system into a more fluent language. This improves the overall system performance.
III Preliminaries
In this section, we introduce the foundational concepts and notation used to describe our two-step neural processing system for EEG-based sentence generation. The system consists of two sequential modules: an encoder for EEG signal processing and a GenAI model to synthesize sentences.
III-A EEG Signal Encoder
Let $\mathbf{X}\in\mathbb{R}^{T\times C}$ represent the EEG input where $T$ denotes the number of time steps and $C$ the number of channels. The encoder function, denoted as $f_{\text{enc}}$, transforms $\mathbf{X}$ into a probability vector $\mathbf{p}\in\mathbb{R}^{26}$, with each component $p_{i}$ of $\mathbf{p}$ representing the probability of the corresponding letter $i$ from the English alphabet:
$\mathbf{p}=f_{\text{enc}}(\mathbf{X};\theta_{\text{enc}})$ (1)
where $\theta_{\text{enc}}$ are the parameters of the encoder. The encoder utilizes a neural network architecture optimized for temporal and spatial features inherent in EEG data.
III-B GenAI Model for Synthesizing Sentences
Upon generating the letter probabilities $\mathbf{p}$, these are inputted into a GenAI model, $g_{\text{trans}}$, which outputs a coherent sentence $\mathbf{y}$. This model is based on a pretrained LLM that has been fine-tuned for the task of converting sequences of letter probabilities into grammatically correct and contextually relevant sentences:
$\mathbf{y}=g_{\text{trans}}(\mathbf{p};\theta_{\text{trans}})$ (2)
where $\theta_{\text{trans}}$ represents the parameters of the translation model. The model leverages advanced natural language processing techniques to ensure semantic coherence and syntactic accuracy in the generated sentences.
These components are integrated to form a robust system for interpreting EEG signals and producing textual outputs, aiming to bridge the gap between neural activities and language expression.
IV Methodology
IV-A Curriculum-based Neural Spelling Framework
This section introduces the Curriculum-based Neural Spelling (CNS) Framework, a novel approach to enhancing the spelling accuracy from neural signals using a two-stage model. Initially, the model utilizes a Convolutional EEG Encoder to transform raw EEG data into a letter classification schema employing minimal neural samples and a tailored set of character categories. In the second stage, we integrate Curriculum Learning with large language models (LLMs) within a sequence-to-sequence framework to generate fluent sentences despite inherent noise and variability in EEG signals. This combination aims to leverage the strengths of advanced neural network architectures and sophisticated natural language processing techniques to overcome the challenges posed by low signal-to-noise ratios and the complex nature of EEG-based letter decoding. The architecture and its components are depicted in Fig. 2, and this comprehensive setup promises significant advancements in brain-to-text communication technologies.
IV-A1 Stage 1: Neural Letter Classifier
Convolutional EEG Encoder
The Convolutional EEG Encoder is designed as a two-layer Convolutional Neural Network (CNN), which is adept at processing EEG data for the purpose of embedding generation, as shown in Fig. 2D. This encoder operates in conjunction with CL techniques to enhance the discriminative power of the embeddings. The trajectory data is processed by a pre-trained ResNet18 [48] model, which serves as a trajectory encoder. The core of our learning strategy is based on minimizing the contrastive learning loss ($loss_{CL}$), which is calculated as follows:
The loss function is defined as:
$loss_{CL}=1-\cos(\theta_{EEG},\theta_{Traj})^{2}$ (3)
where $\theta_{EEG}$ and $\theta_{Traj}$ represent the embedding vectors from EEG data and trajectory data, respectively.
Classification Head
The letter classification component is an integral part of the CNS Framework. It includes a fully connected linear layer that maps its input to a 26-node output layer, where each node represents a letter of the English alphabet. After this. This arrangement enables the system to predict letters based on the processed EEG data. A softmax function is applied to the output layer to obtain a probability distribution over the alphabet, as shown in Fig. 2E, which is critical for the system’s accuracy in spelling prediction. The classifier utilizes the cross-entropy loss to measure the discrepancy between the predicted probabilities and the actual distribution of the target letters. The crossentrypy loss ($loss_{CE}$) function is defined as follows:
$loss_{CE}=-\sum_{c=1}^{M}y_{o,c}\log(p_{o,c})$ (4)
where $M$ is the number of classes (letters), $y_{o,c}$ is a binary indicator (0 or 1) if class label $c$ is the correct classification for observation $o$, and $p_{o,c}$ is the predicted probability of observation $o$ being of class $c$. This loss function effectively guides the learning process by penalizing deviations from the true label distribution, thus enhancing the system’s ability to generate accurate and reliable spelling predictions from EEG signals.
In training phase, the finial loss ($loss_{total}$) is calculated by:
$loss_{total}=0.35\times loss_{CE}+0.65\times loss_{CL}$ (5)
IV-A2 Stage 2: Curriculum Learning for Language Model Integration in Fluent Sentence Generation
The inherently low SNR of EEG signals combined with the rigorous demands of experimental protocols exacerbates the challenges of prediction biases and data scarcity in letter-level classification. These limitations significantly impede the capacity of classifiers to decode fluent sentences. To mitigate these challenges, we propose the integration of LLMs within a sequence-to-sequence (seq2seq) framework aimed at enhancing sentence fluency and readability.
Traditional NLP models often struggle with inputs characterized by significant letter-level noise due to their reliance on sub-word-level tokenization and training predominantly on well-curated texts. As depicted in Table I, without a nuanced understanding of the behaviors specific to subject-based letter classifiers, models like ChatGPT4 can generate sentences that diverge significantly from the intended content, typically relying on generalized knowledge of language structure and common phrases.
TABLE I: Challenges in Letter-Level Noise Denoising with ChatGPT 4
Ground Truth: This civilization, known as the ’Ayar,’ possesses an advanced understanding of nature and the power to manipulate the very fabric of reality.
Prompt: You are an expert on letter denoising of sentences. The letters from the following sentence are corrupted with noise. Can you guess and output the correct sentence? {decoder_output}
Input: tvmg jijifizmhiog wnewn ms tve awmt pesowogeo in iqeanjqv onzerstmgvmng oz nmmxrw mgz txw peeer to mmgipxfamw mxe eefw zavfij ov feifatw
Prediction: Time is the most precious thing in life, understanding it makes the effort to simplify the world of survival or benefit.
Curriculum Learning Approach
To address these constraints, we introduce a curriculum learning strategy tailored for pretrained LLMs. This strategy is designed to adapt the models to the latent distributions specific to subject-based handwriting patterns. By progressively increasing the relevance to the domain and the complexity of the tasks during training, our curriculum learning approach methodically fine-tunes LLMs. This targeted training enables the models to achieve better alignment with the idiosyncratic characteristics of subject-specific letter decoders, thereby enhancing their capacity to handle noisy inputs and improving overall sentence generation:
  1. 1.
    Initial Phase: Start with tasks involving high-frequency, low-noise samples to establish baseline language structures.
  2. 2.
    Intermediate Phase: Gradually introduce more complex and noisier data, increasing exposure to real-world variability.
  3. 3.
    Advanced Phase: Focus on fine-tuning with the highest noise levels and the most challenging samples to ensure robustness and fluency under the most adverse conditions.
This structured approach not only streamlines the adaptation process but also significantly boosts the model’s ability to generalize from noisy, imperfect inputs to coherent and contextually accurate text outputs.
Probabilistic Character Sampling from the Neural Letter Classifiers
To evaluate the efficacy of our curriculum-based LLMs, we analyze the letter category distributions, $C_{i,\omega}$, for letter $\omega$ in sample $i$. These distributions are obtained from the softmax-normalized outputs of the classification head. We compute the aggregate Top-$K$ distributions, $C^{*}_{\omega}$, as follows:
$C^{*}_{\omega}=\frac{\sum_{i=1}^{N}C_{i,\omega}}{N_{\omega}}$ (6)
where $N_{\omega}$ denotes the count of samples for letter $\omega$. From this, the letter $\omega$ is then probabilistically sampled based on its presence within the Top-$K$ distribution. For instance, consider the ground truth (GT) sentence, ”As he chases down rogue time travelers…”. If the aggregated probability for ’A’ at the start, $C^{*}_{A}$, is 0.2 and for ’R’, $C^{*}_{R}$, is 0.6, the classifier might predict ”Rs” in place of ”As”.
Implementation of Noise Adaptation
In this phase, we incorporate the noisy predictions from the Neural Letter Classifier (NLC) into the LLM’s training regimen. Specifically, we replace selected characters in the training texts with their corresponding noisy variants from the Top-$K$ output distributions. The LLM is then fine-tuned to de-noise this altered text. This structured noise adaptation process is designed to enhance the model’s robustness, enabling it to effectively manage and correct noisy inputs during inference.
Curriculum Arrangement
In our curriculum learning method, training examples $\mathbf{D}$ are arranged according to a complexity score $\mathbf{C}$, in a multi-stage setting $\{T_{i}:i=1,2,\cdots,N\}$. In each stage $T_{i}$, a percentage $c\in\mathbf{C}$ of the letters from the input sentences will be permuted by the letter prediction distribution of the letter-classifier model, with a linear increase in the complexity score $c$.
Supervised Fine-tuning
We employ bi-directional attention during training for this seq2seq generation task, as shown in Fig. 2(F). The fine-tuning objective combines the denoising training objective with the causal language modeling objective [45], facilitating the model’s use of previously generated tokens to guide the prediction of subsequent tokens, thereby reducing the influence of noisy letter-level inputs.
$\mathcal{L}_{LMft}=-\sum_{t=1}^{T}\log P(y_{t}\mid\mathbf{x},y_{<t})$ (7)
where $y_{<t}$ represents the previously generated tokens, and $\mathbf{x}$ is the noisy input sentence. In the subsequent Section V-F, we will explore the performance benefits of the fine-tuned LLMs pretrained model utilizing our curriculum learning method in comparison to direct and non-fine-tuning approaches. Additionally, we will examine how variations in model size influence performance metrics.
V Experiments and Results
V-A Participants
We recruited thirty-two right-handed, healthy individuals (P1–P32), all native English speakers from Australia, comprising 16 males and 16 females with an average age of $M_{\text{age}}=24.94\pm 0.29$ years. Each participant had normal or corrected-to-normal vision and reported no history of mental health disorders. 4 subjects are removed from the dataset due to their misunderstanding of the experiment design. Ethical approval was granted by the University of Technology Sydney’s Ethics Committee (Approval No. UTS HREC REF: ETH23-8036). Informed consent was obtained in writing from each participant prior to the experiment.
Figure 3: Model Comparisons: Top 1, Top 3, and Top 5 Accuracy of Different Models. Orange color comparisons shows the difference between w/CL and w/o CL conditions. Asterisks (***) indicate significant results ($\textit{p}~{}<~{}0.001$).
V-B Instrumentation
Handwriting movements and EEG signals were recorded simultaneously at sampling rates of 60 Hz and 1000 Hz, respectively. Handwriting trajectories were captured using a custom-developed application built on the PsychoPy platform, integrated with a Wacom Intuos Pro Medium tablet and Wacom Pro Pen 2 stylus featuring 8192 levels of pressure sensitivity. This application was designed to log critical events and features of each handwriting motion, including timestamps, $x$- and $y$-coordinates, rotation angles, force, and pen state codes for pen-down (contact with the surface), pen-move, and pen-up (lifting from the surface) for each digitalized point along the trajectory.
EEG signals were recorded using a 64-channel Neuroscan amplifier (Curry 9), with electrode placements following the 10–20 international system [49]. A ground electrode was referenced to maintain signal consistency.
Synchronization of handwriting trajectories with EEG signals was achieved by embedding time markers of key events within the EEG stream. Specifically, each first pen-down event of symbols written within each block on the tablet was marked for temporal alignment, as illustrated in Fig. 2(B).
V-C Experimental Design
The experiment was conducted in a sound-attenuated room. Participants were seated comfortably at a desk, positioned approximately 35 cm from a tablet, which was placed in an optimal location on the desk for each individual.
The task followed an event-related design, with extended breaks provided after every 10 trials. During these breaks, participants could rest for an unlimited duration, initiating the next trial by pressing the space key. Each trial, as illustrated in Fig. 2A, began with a fixation cross displayed for 1000 ms with a random jitter of $\pm$ 500 ms. Subsequently, a letter was presented for 2000 ms, cueing the character participants were to write in the handwriting task. During this phase, participants used a stylus to write the letter slowly within a 3000 ms time frame. A 2000 ms relaxation interval followed each handwriting task.
The sequence of the 26 letters was randomized for each participant, with each letter repeated 25 times, resulting in a total of 650 trials $(25\times 26)$ per participant. The entire session lasted approximately 2 hours.
Figure 4: Top 1 Accuracy of CNN w/ CL Model Across Different ROIs and Frequency Bands.
V-D Data Processing
Figure 5: (A): The KDE distributions among only neural letter classifier (NLC), only LLMs (baseline), and LLMs after NLC. (B): The linear regression between the CER and WER scores of NLC ($S_{NLC}$) and LLMs ($S_{LLMs}$).
EEG Preprocessing
EEG data preprocessing was performed using MNE (version 1.6.0) [50] in Python 3.10.13. The machine learning pipeline began with a bandpass filter between 1 and 70 Hz to reduce general noise. This was followed by notch filtering at 50, 100, and 150 Hz to mitigate interference from power lines. Next, epochs were created, defining time intervals from -1 s to 3 s around each event of the first pen-down in each trial. Independent Component Analysis (ICA) was applied with the auto-reject method [51] to remove artifacts from the data. Finally, the EEG data was re-referenced by averaging and baseline correction was applied to ensure consistency across channels.
EEG Feature Extraction
Power Spectral Density (PSD) was extracted as the primary feature for model prediction. To compute PSD, an Fast Fourier Transform (FFT) was used, capturing frequencies between 1 Hz and 70 Hz, as shown in Fig. 2C. The PSD for a given signal $x(t)$ was calculated using the formula:
$\text{PSD}(f)=\frac{1}{N}\left|\sum_{t=0}^{N-1}x(t)e^{-i2\pi ft/N}\right|^{2}$ (8)
where $f$ represents the frequency, and $N$ is the total number of time points in the signal. This approach provided a detailed frequency profile of the EEG signals, essential for downstream analyses.
Trajectory Processing
Trajectory data was processed by applying min-max normalization to the $x(t)$ and $y(t)$ coordinates, which were subsequently mapped onto a 28 x 28 pixel grid. To represent temporal progression, the intensity values along the trajectory were scaled from 50 to 255, illustrating the passage of time, as shown in Fig. 2B. This approach allowed the temporal aspects of each handwriting movement to be visually represented in the spatial grid format.
V-E Stage 1: Neural Letter Classifier
V-E1 Model Performance Comparison
In this section, we evaluate the efficacy of different neural network architectures for EEG-based encoding by comparing a CNN with both Long Short-Term Memory (LSTM) networks and Transformer-based models, each equipped with and without CL module. The comparative analysis is rooted in statistical testing and performance metrics across three classification accuracies: Top 1, Top 3, and Top 5.
TABLE II: Main Decoding Results
arch Curriculum BLEU (%) ROUGE (%) Error Rates (%)
1 2 3 4 1 2 L Lsum CER WER
Bart-base no 25.74 10.77 5.02 2.68 22.78 3.80 18.51 18.50 92.61 120.32
yes 27.55 11.46 5.64 3.17 24.07 4.04 19.51 19.50 89.63 113.22
Bart-large no 58.52 47.86 40.91 35.76 58.18 41.13 57.11 57.10 44.96 53.54
yes 64.89 55.56 49.27 44.42 64.63 49.91 63.72 63.71 38.94 46.68
TABLE III: Ablation on the impact of space separation between words during neural spelling
arch Space BLEU (%) ROUGE (%) Error Rates (%)
1 2 3 4 1 2 L Lsum CER WER
Bart-base no 40.14 25.47 18.24 13.97 35.24 15.12 32.00 31.97 66.46 83.10
yes 45.46 32.01 24.92 20.37 41.15 21.96 38.30 38.27 60.89 76.22
Bart-large no 58.52 47.86 40.91 35.76 58.18 41.13 57.11 57.10 44.96 53.54
yes 64.89 55.56 49.27 44.42 64.63 49.91 63.72 63.71 38.94 46.68
TABLE IV: Ablation Study on the impact of letter classification sampling space
Model Arch Sample Space BLEU (%) ROUGE (%) Error Rates (%)
1 2 3 4 1 2 L Lsum CER WER
Bart-base 3 58.52 47.86 40.91 35.76 58.18 41.13 57.11 57.10 44.96 53.54
7 51.79 39.32 31.89 26.80 49.31 30.32 47.60 47.60 53.69 64.31
11 45.98 32.34 24.72 19.83 42.78 22.54 40.56 40.54 60.31 72.27
15 44.62 30.68 23.06 18.29 40.97 20.65 38.63 38.61 62.08 74.78
19 43.76 29.65 22.10 17.42 40.01 19.59 37.53 37.52 63.12 76.09
23 43.48 29.31 21.74 17.07 39.65 19.22 37.11 37.09 63.46 76.55
26 48.14 34.61 27.03 22.10 44.44 24.68 42.14 42.14 58.78 71.18
Bart-large 3 64.89 55.56 49.27 44.42 64.63 49.91 63.72 63.71 38.94 46.68
7 55.22 43.30 35.94 30.73 52.74 34.52 51.12 51.13 50.72 60.95
11 51.25 38.41 30.86 25.76 48.06 28.87 46.06 46.06 55.41 66.70
15 49.28 36.03 28.44 23.44 45.73 26.16 43.52 43.53 57.64 69.62
19 48.44 35.01 27.43 22.48 44.94 25.19 42.67 42.67 58.31 70.47
23 47.94 34.37 26.78 21.85 44.30 24.48 41.98 41.97 59.13 71.60
26 48.13 34.66 27.11 22.20 44.44 24.74 42.12 42.13 58.92 71.36
The CNN augmented with the CL module significantly outperformed the other models. Specifically, the CNN w/ CL achieved a Top 1 accuracy of $33.10\%\pm 11.54\%$, a Top 3 accuracy of $57.77\%\pm 12.89\%$, and a Top 5 accuracy of $71.46\%\pm 11.11\%$. Statistical analyses, including pairwise t-tests, revealed that the CNN with CL model was superior to the LSTM w/ CL (t(334) = 9.505, $\textit{p}~{}<~{}0.001$) and the Transformer w/ CL (t(334) = 12.028, $\textit{p}~{}<~{}0.001$). All tests are corrected by Benjamini-Hochberg False Discovery Rate (FDR_BH) to avoid multi-comparison issues, ensuring robustness in reporting statistically significant differences.
Fig. 3 delineates the comparative accuracy of these models across the Top 1, Top 3, and Top 5 metrics, where asterisks denote statistical significance at the $\textit{p}~{}<~{}0.001$ level.
V-E2 Model Performance Across Different ROIs and Frequencies
This section investigates the impact of different ROIs and frequency bands on the Top 1 accuracy of a CNN model augmented with CL. The analysis focuses on the PSD extracted from specific bands and ROIs to determine their relative importance in performance metrics.
Fig. 4 illustrates the results of this comparison. Utilizing all features from all bands and ROIs yields the highest Top 1 accuracy, indicating the advantage of a comprehensive feature set. The Gamma band stands out among the frequency bands, showing the largest performance improvement, underscoring its significance in EEG-based models. Regarding the ROIs, the PFC achieves the first-highest Top 1 accuracy, followed by the whole frontal cortex. Subsequent analyses show that the temporal and sensorimotor cortices rank lower, yet they contribute to overall model accuracy.
V-F Stage 2: Curriculum-based Large Language Model
V-F1 Sentence Dataset
For Stage 2 verification, we utilize the Story-Plots-1.3k dataset from QuasarResearch, an expansion of the Neural-Story-v1 from NeuralNovel. This dataset is specifically designed for applications in creative text generation and narrative analysis, comprising 1,320 unique story plots. All stories are included in our testing set. The dataset is hosted on Hugging Face, available at https://huggingface.co/datasets/QuasarResearch/story-plots-1.3k.
V-F2 Sentence Decoding Performance
Table II presents the results of sentence decoding from the encoder, utilizing the Top-k predicted distribution as input for the LLM decoder (where $k=3$). We evaluate the decoding performance across two different model sizes of the BART model [45], namely ‘large‘ and ‘base‘, with the trainable parameters equal to 406M and 140M, respectively. For both model sizes, our proposed curriculum learning method significantly enhances decoding performance across all evaluated metrics. Specifically, using the pretrained BART-large model, our method achieves a BLEU-4 score of $44.42\%\pm 2.2\%$ and a ROUGE-L score of $63.71\%\pm 0.86\%$.
Additionally, Fig. 5(A) compares the performances among only neural letter classifier, only LLMs (baseline), and LLMs after neural letter classifier, show the LLMs can improve the performance for this spelling-based language decoding. This observation underscores the potential of LLMs to enhance the language decoding task, suggesting that improvements in letter classification directly contribute to better sentence reconstruction outcomes.
Fig. 5(B) revealing a positive linear relationship between the decoder’s performance on the sentence decoding task and the neural letter classifier’s performance. It shows that LLMs after neural letter classifier can significantly improve in word level.
V-F3 Visualization of Sentence Decoding Results
TABLE V: Main Decoding Results
Type Content
Classifier Output rn thp unltrgfving wflyernesf kl the amaztn fafnforelt i sroup os rplgrfchprs lgd bu da rtpria srnvhpz uncowgr r grydgn cicilrzaorkn dggp wrohfn thg kunsle
Ground Truth In the unforgiving wilderness of the Amazon rainforest, a group of researchers led by Dr. Amelia Sanchez uncover a hidden civilization deep within the jungle.
LM Decoder In the unforgiving winds of the Amazon rainforest, a group of survivors led by Dr. Anya Singh uncover a hidden civilization that harnessed the power of nature’s healing properties.
Classifier Output al hef creitikns bpuowg fncrgrsingry xtoyfrf fhe frups thg gthfcrl dfrgtma of trnfxulrtrnj xpooref orlts ind the ctnsequencel ol blyfrrng zhg rfnpf bgtwepn rparity rnd srbrication
Ground Truth As her creations become increasingly powerful, she faces the ethical dilemma of manipulating people’s emotions and the consequences of blurring the lines between art and reality.
LM Decoder As her creations become increasingly popular, she faces the ethical dilemma of manipulating people’s pasts and the consequences of blurring the lines between reality and fabrication.
Classifier Output tfiver r cybtrg detpctivp mith the ibilrty tt rewind trmg fnceszfgazgs tempoaar crrtef rn r futurfszrv tetroooril
Ground Truth Oliver, a cyborg detective with the ability to rewind time, investigates temporal crimes in a futuristic metropolis.
LM Decoder Thomas, a cybernetic detective with the ability to rewind time, investigates temporal anomalies in a futuristic metropolis.
Classifier Output as ag chrfef yown rogue trmp trrvglgrl and ynaavpff uknsxifauiel ag unvocprf i lakcking zruth rbout hff omn crpaoitn rnd fts oiel oo thg very frbrru ks tfte skauing hrt zt uhtklg setwgen lgavrnj duszice tr embrrcrng hrl taue oufxole even rs io mgrns rloering zhe ctyrse of hiltkay
Ground Truth As he chases down rogue time travelers and witnesses unimaginable events, he uncovers a shocking truth about his own creation and its potential to shape the future of time, forcing him to choose between preserving history or allowing it to evolve naturally.
LM Decoder As he chases down rogue time travelers and unravels conspiracies, he uncovers a shocking truth about his own creation and its ties to the very fabric of time, forcing him to choose between serving justice or embracing his true purpose, even if it means altering the course of history.
Classifier Output fn a xoltaotcrlyotrc waftefiny krf i rgsoyruesur lvavenger ltumbles uxtn r soagkzten rfvhrvp cknzainrng loabfkken knomlpdgg rbtut thp wkrld bestap tae ctrrrolg
Ground Truth In a post-apocalyptic wasteland, Kai, a resourceful scavenger, stumbles upon a forgotten archive containing forbidden knowledge about the world before the collapse.
LM Decoder In a post-apocalyptic world, Kai, a resourceful scavenger, stumbles upon a forgotten archive containing forgotten knowledge about a legendary artifact.
Classifier Output kn the dusty tarfls ks thg grrnd vrnutn srgp r nowrdrv artilt wioa a paratbrush rf hea gurdg lharpl ftkrfgs wrza zeohya r wffe old ckykte wgo underftrnyl the rgyohmf tl thg kesgrt
Ground Truth On the dusty trails of the Grand Canyon, Sage, a nomadic artist with a paintbrush as her guide, shares stories with Zephyr, a wise old coyote who understands the rhythms of the desert.
LM Decoder On the dusty plains of the Grand Canyon, Sage, a nomadic artist with a brush as her guide, shares stories with Whispering, a wise old Cherokee woman who understands the language of the land.
Note: All letters in the displayed results have been automatically converted from uppercase to lowercase for visualization.
For qualitative analysis, we visualize the decoding results of our method in Table V. The input samples were taken from subject P11, which exhibits the highest performance in letter classification. Notably, some phrases from the initial decoding are already remarkably similar to the ground truth. For instance, ’rn thp unltrgfving…’ closely approximates ’in the unforgiving…’. Our LLM decoder capitalizes on the word-transitional distribution, facilitating improvements in sentence grammar and word-level corrections. As a result, the decoded sentences closely align with the ground truth, demonstrating the efficacy of our approach in refining linguistic outputs.
Figure 6: Significance of EEG spectral variations across different regions and frequencies, analyzed using F-statistics (F(25,675)) and corrected for multiple comparisons with the FDR_BH approach (colors indicate $\textit{p}<0.05$).
V-G Abulation Study
V-G1 Ablation on the effect of letter sampling range
In our analysis, we observe a decline in performance across all metrics when sampling from a larger prediction range. This outcome illustrates a trade-off between letter-level prediction accuracy and input diversity. To balance these factors effectively, we opt to utilize the top-3 prediction results for each letter to sample the letters spelled by the user, as detailed in Table IV. This approach optimizes our model’s accuracy while maintaining a reasonable level of diversity in the input data.
V-G2 Ablation on the Effect of Space Between Words During Neural Spelling
Results are presented in Table III. In conclusion, while incorporating spaces between words introduces an additional step for the user, it aids the model in more effectively denoising the input. However, the overall results demonstrate that our model is capable of denoising neural spelling text effectively even when spaces are not included in the input. This underscores the robustness of our approach in handling continuous streams of characters without explicit segmentation.
VI Discussion
This study presents a Curriculum-based Neural Spelling Framework comprising a Convolutional EEG Encoder and a CL module, demonstrating the feasibility of recognizing handwritten letters using non-invasive technologies. Furthermore, the system highlights the potential for real-world applications by integrating LLMs to generate free-form sentences based on spelling-based designs.
Figure 7: Visualization of ITC across different frequencies (1Hz to 70 Hz) during handwriting tasks. The color scale indicates the level of coherence, with warmer colors representing higher coherence.
VI-A Neurophysiological Correlates of Letter Recognition
Figure 8: UMAP visualization of letter-specific neural representations. (A): EEG PSD features with averaged baseline removed to enhance visualization clarity. (B): Handwriting trajectories showing distinct clusters for letters TWY, BFE, ZK, and OC, which suggest similar neurocognitive processing pathways and minor differences in movement patterns. (C): Significant differences between letter pairs A and Z, B and L, B and J, and O and Z, assessed via paired t-tests and corrected for multiple comparisons using the FDR_BH method (colors indicate $\textit{p}<0.05$), highlighting the neurophysiological distinctions in letter processing.
Handwriting involves complex neural coordination, engaging multiple brain regions. Figure 8 illustrates how specific letters correspond to neural activation (A) and handwriting trajectory (B). Some letters, such as TWY, BFE, ZK, and OC, show similar patterns in both neural and trajectory patterns. The small distance of these letters, suggests a shared representation in neural processing spaces.
Significant neurophysiological variations are highlighted in Fig. 6, where the F-statistics reveal prominent differences across cerebral regions at specific frequency bands. In particular, the alpha band shows substantial activity in the occipital and parietal lobes ($\textit{p}<0.05$), underscoring their roles in visual processing and spatial integration [52, 53], essential for interpreting letter shapes and trajectories.
As demonstrated in Fig. 8(C), the neural pattern differences between letters exhibit complex and diverse styles across different brain regions and frequency bands ($\textit{p}<0.05$). Specifically, the differences between letters A and Z are predominantly observed in the Gamma band within the Frontal and Parietal cortices ($\textit{p}<0.05$). In contrast, the patterns for B and L are distinctly marked by Alpha band activity in the Parietal and Central cortices ($\textit{p}<0.05$). Interestingly, the neural patterns for B and J are similar to those of B and L but feature less Gamma band activity in the Frontal cortex ($\textit{p}<0.05$). Moreover, the patterns for O and Z resemble those between A and Z, albeit with a reduced difference in Gamma activity ($\textit{p}<0.05$). These observations highlight the unique neurophysiological pathways for each letter pair, reflecting varied motor and cognitive demands, and suggesting differences in cognitive processing levels.
The frontal cortex demonstrates increased gamma-band activity, reflecting its integrative function across sensory modalities and its pivotal role in higher cognitive processes, such as memory and decision-making [54, 55, 56, 57], which are crucial for language processing and handwriting execution.
Moreover, the prefrontal cortex (PFC) emerges as a critical node, facilitating the integration of multi-modal information and mediating complex cognitive functions including executive control and working memory [58, 59]. Similarly, temporal cortex activations are closely tied to language processing, with gamma activity playing a significant role in the neural decoding of language elements [7, 5].
Additionally, the sensorimotor cortex’s involvement aligns with its role in governing motor control and sensory processing, fundamental to handwriting [60]. The broad gamma activity across these regions suggests a high-level synchronization of cortical activity, facilitating coherent cognitive representations necessary for complex tasks like letter recognition and differentiation [61, 62, 63].
This study underscores the nuanced interplay of multiple brain regions during letter recognition and handwriting, paralleling findings from related fMRI studies [64]. The identified patterns across the frontal, parietal, and temporal cortices substantiate their critical roles in the cognitive and perceptual foundations of language and handwriting tasks. Utilizing all ROIs and frequency bands yields the best performance, as demonstrated in Fig. 4.
VI-B Inter-Trial Coherence in Delta Band During Handwriting Tasks
Inter-Trial Coherence (ITC) measures phase synchronization across EEG trials, providing insights into consistent neural responses and cortical synchrony during event-related cognitive tasks [65]
$ITC(f)=\left|\frac{1}{N}\sum_{k=1}^{N}e^{i\phi_{k}(f)}\right|$ (9)
where $N$ is the number of trials, $f$ represents the frequency, and $\phi_{k}(f)$ is the instantaneous phase of the $k$-th trial at frequency $f$.
Our analysis, during handwriting tasks, particularly underscores the Delta frequency band, which exhibits significant coherence during the initial phase of action execution (0 to 1 second), as depicted in Fig. 7. This heightened coherence suggests a robust engagement of neural circuits associated with low-frequency oscillations, crucial for the timing and coordination of motor movements.
Studies have indicated the involvement of Delta and Theta bands in motor tasks and cognitive processing. For instance, increased ITC in these bands has been associated with task switching and inhibitory control in BCI paradigms [66]. Similar patterns of coherence in Delta-Theta ranges have been observed during motor performance improvements in stroke rehabilitation, reflecting neural reorganization and recovery [67]. The periodic auditory stimulation research shows a synchronization peak at 2 Hz in the Delta band, linking sensory-motor time coupling, which may parallel the neural dynamics observed during the rhythmic movements in handwriting [68].
Moreover, the Alpha band, while showing lower coherence compared to the Delta band during the initial phase of handwriting, plays a significant role in broader cognitive and motor control contexts. Increased Alpha ITC has been associated with enhanced visuomotor performance and is particularly evident in tasks requiring visual coordination, such as visuomotor tracking [69]. Furthermore, variability in Alpha ITC has been linked to cognitive control efficiency within the frontoparietal network [70], an aspect crucial for complex task execution.
This differential engagement of frequency bands underscores the complexity of brain dynamics during fine motor tasks such as handwriting. While Delta, Theta, and Alpha oscillations may not yield the greatest contributions in NLC applications, their primary roles in initiating and coordinating motor output, coupled with their subsequent involvement in cognitive processing, illustrate a layered neural architecture that supports both the execution and cognitive integration of handwriting tasks. Although these bands do not dominate the neural decoding efforts, their influence is nonetheless crucial for the holistic understanding of motor control mechanisms.
VII Conclusions
This study introduces a Curriculum-based Neural Spelling Framework (CNS) that leverages the advanced capabilities of GenAI to enhance spell-based neural language decoding tasks. Our approach is distinct in integrating a CNN with a curriculum-driven LLM, promoting an innovative hybrid method in the domain of BCIs. The framework’s effectiveness is demonstrated through its application to EEG-based handwriting of all 26 letters, a novel endeavor in the field. The CNS framework notably achieves exemplary top-k accuracy across all subjects, underscoring the robustness of the EEG encoding model. Furthermore, our curriculum supervised fine-tuning method significantly advances the state of the art by enabling the LLM to effectively learn subject-specific letter transition patterns. This methodological innovation not only enhances sentence synthesis quality but also sets a new benchmark for assistive communication technologies. By seamlessly merging non-invasive EEG with GenAI, this study not only addresses the immediate needs of individuals with diverse physical abilities but also sets the stage for future explorations in sophisticated, accessible communication solutions. As we continue to refine these integrations, the potential to expand the capabilities of BCIs and improve the quality of life for users worldwide remains vast and inspiring.
VIII Limitations
Despite its notable strengths, the proposed framework exhibits several limitations that warrant further investigation. Firstly, there is a need to collect and validate online sentence data to ascertain the framework’s efficacy in real-world applications. Secondly, the current task framework is primarily based on within-subject analyses, limiting its generalizability across different subjects. The capacity for cross-subject transfer learning remains underexplored and requires significant enhancement to ensure broader applicability. Thirdly, the dataset size for individual subjects is relatively small, which constrains the training of a more extensive and robust brain encoder necessary for a high-performing letter classifier. Future work should focus on collecting more comprehensive data sets and developing a robust, transferable model that can operate effectively across subjects in real-world, online applications.